40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 28 23:33:38 2020, gautam, Update, BHD, Some more hardware changes 
    Reply  Sat Aug 29 22:46:29 2020, gautam, Update, BHD, New homodyne-phase control electronics zetaDrive.pdftrekTFs.pdf
       Reply  Wed Sep 2 00:49:47 2020, gautam, Update, BHD, Some notes about homodyne phase heterodyneMICH.pdfunwrappedPhase.pdfunwrappedPhase_zoom.pdfphaseNoisePSD.pdf
          Reply  Thu Sep 3 15:55:04 2020, gautam, Update, BHD, Phase drift between LO and IFO after fiber replacement LOphaseDrift.pdfphaseDrift_tempCorr.png
Message ID: 15555     Entry time: Thu Sep 3 15:55:04 2020     In reply to: 15553
Author: gautam 
Type: Update 
Category: BHD 
Subject: Phase drift between LO and IFO after fiber replacement 

Summary:

After replacement of the fiber delivering the LO beam to the airBHD setup (some photos here), I repeated the measurement outlined here. There may be some improvement, but overall, conclusions don't change much.

Details:

The main addition I made was to implement a digital phase tracker servo (a la ALS), to make sure my arctan2 usage wasn't completely bonkers (the CDS block can be deleted later, or maybe it's useful to keep it, we will see). I didn't measure it today, but the UGF of said servo should be >100 Hz so the attached spectrum should be valid below that (loop has not been done, so above the UGF, the control signal isn't a valid representative of the free running noise). Attachment #1 shows the result. The 1 Hz and 3 Hz suspension resonances are well resolved. Anyways, what this means is that the earlier result was not crazy. I don't know what to make of the high frequency lines, but my guess is that they are electronic pickup from the Sorensens - I'm using clip-mini-grabbers to digitize these signals, and other electronics in that rack (e.g. ALS signals) also show these lines.

It is pretty easy to keep the simple Michelson locked for several minutes. Attachment #2 shows the phase-tracker servo output over several minutes. The y-axis units are degrees. If this is to be believed, the relative phase between the two fields is drifting by 12um ove an hour. This is significantly lower than my previous measurement, while the noise in the ~0.5-10 Hz band is similar, so maybe the shorter fiber patch cable did some good?

I think there is also correlation between the PSL table temperature, but of course, the evidence is weak, and there are certainly other effects at play. At first, I thought the abrupt jumps are artefacts, but they don't actually represent jumps >360 degrees over successive samples, so maybe they are indicative of some real jump in the relative phase? Either fiber slippage or TT suspension jumps? I'll double check with the offline data to make sure it's not some artefact of the phase tracker servo. If you disagree with these conclusions and think there is some meaurement/analysis/interpretation error, I'd love to hear about it.

Next steps:

  1. Budget the offline inferred phase noise spectrum, overlay a seismic noise model, to see if we can disentangle the contributions from the suspensions and that from the LO fiber.
  2. I'll see if I can setup an LO pickup with some RF sidebands on it in parallel to this setup so we can try some of the ideas discussed on the call this week. There are several beams available, but the question is whether I can get this into a fiber without 1 week of optical layout work.

I have left the heterodyne electronics setup at the LSC rack, but it is not powered (because there are some exposed wires). Please leave it as is.

Attachment 1: LOphaseDrift.pdf  27 kB  Uploaded Thu Sep 3 17:12:11 2020  | Hide | Hide all
LOphaseDrift.pdf
Attachment 2: phaseDrift_tempCorr.png  104 kB  Uploaded Thu Sep 3 19:43:04 2020  | Hide | Hide all
phaseDrift_tempCorr.png
ELOG V3.1.3-